hyperbula 4y^2-2x^2=12 open to the left or right?
Answer by lwsshak3(11628)

hyperbola 4y^2-2x^2=12 open to the left or right?
4y^2-2x^2=12
divide by 12
y^2/3-x^2/6=1
This is an equation of a hyperbola with a vertical transverse axis (y-term comes first), so hyperbola opens up and down.